OpenStudy (phebe):

MEDALSS!!! & fann A 12ft by 15ft rectangular swimming pool has a 3ft wide no slip surface around it. What is the outer perimeter of the no-slip surface? A- 78ft B-78ft^2 C-198ft D-198ft^2

OpenStudy (phebe):

can u help me with this

OpenStudy (anonymous):

12+3+3 = 18,* 2 sides = 36 15+3+3= 21 * 2 sides = 42 add the 4 sides, 36+42 = 78 (A)

OpenStudy (phebe):

thanx can u help me with another one nd are u sure i don't add ^2

OpenStudy (anonymous):

I'm sure. And i'll try!

OpenStudy (phebe):

ok thanx

OpenStudy (phebe):

want me to put it in another post or this one

OpenStudy (anonymous):

You can put it in this

OpenStudy (phebe):

ok hold on k

OpenStudy (phebe):

What is the perimeter of PQR with vertices P(-2,9), Q(7,-3), and R(-2,-3) in the coordinate plane?

OpenStudy (phebe):

A-21units B-25 units C-34units D-36units

OpenStudy (anonymous):

PQ = 15 QR = 9 RP = 12 The perimeter is 15 + 9 + 12 = 36.

OpenStudy (phebe):

thanx at least now i know how to do it, what grade u in?

OpenStudy (phebe):

sorry but can u help me with another

OpenStudy (anonymous):

I'm in 9th And umm... Sure i'll try.

OpenStudy (phebe):

ok kcool nd thanx

OpenStudy (phebe):

you are tiling a kitchen floor that is 10ft. wide by 4yd long. How many square yards of tile do you need?

OpenStudy (phebe):

A-13and 1/3yd^2 B- 13 and 1/2yd^2 C-20yd^2 40yd^2

OpenStudy (phebe):

D-40yd^2

OpenStudy (anonymous):

You will need 13 and 1/3 square yards of tile.

OpenStudy (phebe):

u sure jus askin

OpenStudy (phebe):

not trying to giv u all my work but cna u help me with another one

OpenStudy (phebe):

nvm